# Probing Green’s Function Zeros by Co-tunneling through Mott Insulators

### v1

@@ A.5 Self-consistent selfenergy construction
Assuming a small $||\mathcal{S}^{-1}\tilde{G}^{-1}G_{0}^{-1}||<1$ (self-consistency condition), we can expand the right site as a Neumann series and find:
@@
As last step we multiply by $\mathcal{S}^{-1}$ from right and expand both ,$\mathcal{G}_{\text{H}}$ and $\mathcal{S}^{-1}$, in orders of $\phi$ (up to the highest known order $n$ of $\mathcal{G}_{\text{H}}$ ):
$\mathcal{G}_{\text{H}}$ $\approx\sum_{k=2}^{n}\phi^{k}\mathcal{G}_{\text{H},k}+\mathcal{O}(\phi^{n+1})\text{ ,}$ (64)
$\mathcal{S}^{-1}$ $\approx\sum_{k=0}^{n}\phi^{k}\mathcal{S}^{-1}_{k}+\mathcal{O}(\phi^{n+1})\text{ ,}$ (65)
@@
Within this letter we already derived the GF $G$ in Eq.(34) up to the second order in $\phi$:
@@
the natural decomposition is then:
$\tilde{G}^{-1}$ $=-\left[H^{z}-\omega\right]$ (70)
$\mathcal{G}_{\text{H}}$ $=-\frac{\phi^{2}}{D^{2}}\mathbb{I}$ (71)
@@
$\mathcal{S}^{-1}_{0}$ $=0$ (72)
$\mathcal{S}^{-1}_{1}$ $=0$ (73)
$\mathcal{S}^{-1}_{2}$ $=\frac{D^{2}}{\phi^{2}}\mathbb{I}\text{.}$ (74)
@@
$\Sigma$ $\approx\frac{D^{2}}{\phi^{2}}\left[\omega-H^{z}\right]^{-1}$ (75)
Appendix B Perturbation theory for two auxiliary spinful quantum dots
@@
[…]
@@
$\mathcal{H}^{D}$ =$=\{|0,\uparrow\rangle,|0,\downarrow\rangle,|\uparrow,0\rangle,|\downarrow,0\rangle\}$ (80)
@@
The GFZ-Hamiltonian $H^{\text{z}}$ was already derived in Eq.(55), while $H_{0}$ denotes the bare non-interacting Hamiltonian. The result is valid for $\omega\ll U$, where either $\alpha=0$ or $|\alpha|\geq 2$. For $20<|\alpha|<$ 2 the expansion is close to the singularities in the Hubbard band. Together with Eq.(82) we can further simplify:
@@ Appendix C Charge stability diagram
We compute the stability diagram by determining the global groundstate of the full Hamiltonian $H_{\text{tot}}$ in Eq.(76) for varying weak dot potentials $\mu_{\text{L}},\mu_{\text{R}}$. The groundstate obeys a half-filled Hubbard system for small potentials $\mu_{\text{L}},\mu_{\text{R}}$, while there are either zero $(0,0)$, one $(1,0)$ (dominantly at the left dot), $(0,1)$ (dominantly at the right dot) or two $(1,1)$ particles at the auxiliary dots. We then determine the dot occupations $N_{L}(\mu_{\text{L}},\mu_{\text{R}}),N_{R}(\mu_{\text{L}},\mu_{\text{R}})$. With the onset of the co-tunneling induced coupling between the auxiliary dots, we find that there is a smooth transition between the two localized regimes $(1,0)$ and $(0,1)$. The resulting charge curve $N_{L}(\mu_{\text{L}},\mu_{\text{R}})$ ,$N_{R}(\mu_{\text{L}},\mu_{\text{R}})$ in this regime can be derived from Eq.(90). Let us rewrite:
@@
Where we used that $a_{s}(\omega)$ and $b_{s}(\omega)$ are linear in the frequency $\omega$. Now for the eigenstates we can neglect the overall shift $C_{0}$ and for a fixed $\bar{\mu}$ introduce the shifted potential difference $\tilde{\delta}\mu=\delta\mu+\frac{C_{1}\bar{\mu}+C_{3}}{C_{2}}$:
@@
$|\varepsilon_{-}\rangle$ $=\frac{1}{\sqrt{\left(\chi-\sqrt{\chi^{2}+1}\right)^{2}+1}}\begin{pmatrix}\chi-\sqrt{\chi^{2}+1}\\ 1\end{pmatrix}$ (98)
@@
The occupations on the dots $N_{L}(\bar{\mu},\tilde{\delta}\mu)$,$N_{R}(\bar{\mu},\tilde{\delta}\mu)$ are then given by:
@@
The Full-Width-at-Half-Maximum is then given by $\Delta_{FWHM}=2\sqrt{2^{\frac{2}{3}}-1}\frac{|t_{\text{co}}(\bar{\mu})|}{C_{2}}\approx 1.53\frac{|t_{\text{co}}(\bar{\mu})|}{C_{2}}$. Moreover, we can approximate $C_{2}\approx\frac{1}{2}+\mathcal{O}(\frac{\gamma^{2}}{U^{2}})$ and in the case of a strictly local system-dot couplings $l_{k}\propto\delta_{ik},r_{k^{\prime}}\propto\delta_{jk^{\prime}},i\neq j$, the co-tunneling $t_{\text{co}}$ is also independent of $\bar{\mu}$. In total we find a universal curve with $\Delta_{FWHM}\approx 3.07|t_{\text{co}}|$.
Appendix D Deviations for state-of-the-art quantum dot parameters
Figure 5: Co-tunneling $t_{\text{co}}$: numerical results (ED) vs analytical results (Eq.(93)) for the SSH-Hubbard chain ($N=16$, $w=0.8v$, PBC, $l_{i}=\delta_{i,1}$,$r_{i}=\delta_{i,2}$). In (a): Co-tunneling $t_{\text{co}}$ as function of $\alpha$ and different relative interaction strengths $\phi$, $D=100$ fixed, numerical results(dashed, green) and analytical results(solid,red). In (b): relative error of $t_{\text{co}}$ determined by Eq.(93) compared to the exact result for various $D$,$U$ and fixed $\gamma=30$. In (c): absolute value of $t_{\text{co}}$, same parameters as in (b). As reference: contour lines of constant $\alpha$, $\phi$ and $t_{\text{co}}$ (grey, green,black).
@@ Appendix D Deviations for state-of-the-art quantum dot parameters
Both are sufficient to describe the universal behavior, as a global rescaling of the original parameters does not affect the underlying physics. The previously discussed parameter $\phi$ characterizes the Mottness of the central Hubbard system, while $\alpha$ represents the ratio of $\gamma$ to the well-known exchange coupling $J=\frac{D^{2}}{U}$, which corresponds to the half-bandwidth of the effective low-energy Heisenberg model. In Fig. 5(a), we present the co-tunneling amplitude $t_{\text{co}}$ as function of the coupling ratio $\alpha$ and for various relative interaction strengths $\phi$ in the SSH-Hubbard chain ($w=0.8v$, PBC, $D=(v+w)/2$) with local couplings $l_{i}=\delta_{i,1}$ and $r_{i}=\delta_{i,2}$:
@@
We find that the exact results (dashed, green) show excellent agreement with our analytical approach (solid, red), as given by Eq. (90), already at $\phi\approx 0.2$ and up to relatively large values of $\alpha\approx 2$. As expected, increasing Mottness and weaker lead-system coupling cause the exact results to converge toward the analytical predictions. More surprising, however, is the large value $\alpha$ with a good agreement, which we interpret as effect of the vanishing first-order contributions of the weak lead-system hopping $C$ in Eq.(78). The drastic change for $\phi>0.2$ signals the Mott transition. From Fig. 5(a), the bare parameters $\gamma,U$ and $D$ can be directly extracted. For example, for $\phi=0.2$ and $\alpha=1.5$ let us fix a half-bandwidth $D=100\mu eV$, then we obtain $U=500\mu eV$, $\gamma=30\mu eV$, and $t_{\text{co}}\approx 1\mu eV$.
@@
Appendix E Spin renormalization of the GFZ Hamiltonian
@@
Figure 6: In (a): charge oscillations between aux. dots (initial localization left dot, ED,$U=200$ $N=20$, $\gamma=2$, $v=18$, $w=20$, OBC), with $l_{i}=\delta_{i,1}$,$r_{i}=\delta_{i,2}$ (upper,violet) and $l_{i}=\delta_{i,3}$,$r_{i}=\delta_{i,4}$ (lower,green). Analytical result (dashed) governed by Eq.(90). In (b): staggering for OBC inverting at the boundary close to topological transition $v=w$. Same parameters as in $(a)$.
@@
We comment on the renormalization of the GFZ-Hamiltonian of Eq.(55) by the spin-spin correlation functions. In particular, we focus on the topological features of the underlying SSH model. We consider a finite-size SSH-Hubbard chain, where the non-interacting system shows zero-energy edge modes as a function of the staggered hopping ratio $v/w$. As described by to Eq.(55), the bare hoppings get renormalized significantly by the spin-correlation values. Since the SSH model at large U is effectively a spin chain, its ground state will obey Affleck-Kennedy-Lieb-Tasaki (AKLT) physics [70], with the formation of spin-singlet dimers in correspondence with the stronger bonds. As a consequence, the original imbalance between the non-interacting $v$ and $w$ is further amplified. In Fig.(6)(b) we show the behavior of the renormalized ratio $\frac{\tilde{v}}{\tilde{w}}$ as a function of $v$, for $w=20$ and $U=200$. The grey line shows the behavior of the bare hopping ratio. It is immediate to notice how the renormalized ratio varies much more heavily, passing from strong suppression to strong enhancement. The purple and green curves in the plot refer to different positions along the finite SSH chain, respectively near the boundary and deep in the bulk. Finite-size effect entail a different renormalization in the two cases, and the topological transition, denoted by the red dots, happens at the boundary sites for a lower value of $v/w$ than in the bulk. In the region between the two red dots, the boundary sites are already in the trivial phase, while the bulk is topological. As a consequence, edge state reconstruction ensues [71], with the topological edge zero described in [33] shifting into the bulk.
Appendix F Minimal chain to resolve the spin renormalization
@@
Figure 7: In (a) and (b): small chain of two coupled Hubbard dimers and lead dots at site 1 and 2 $l_{i}=\delta_{i,1}$,$r_{i}=\delta_{i,2}$. (a): Two uncoupled dimers with spin-singlet at bonds 1-2 and 3-4, probed $t_{\text{co}}$ at bond 1-2 is enhanced by the spin correlation $\langle\overrightarrow{S}_{1}\cdot\overrightarrow{S}_{2}\rangle$. (b): spin-singlet at bonds 2-3 and 1-4, probed $t_{\text{co}}$ at bond 1-2 is suppressed by the spin correlation $\langle\overrightarrow{S}_{1}\cdot\overrightarrow{S}_{2}\rangle$. In (c) and (d): expected (normalized) co-tunneling $t_{\text{co}}$ as function of $w$ and $U$ (ED, $N=12$, fixed $U=400$ on the lead dots, $v=100$, $l_{i}=\delta_{i,1}$,$r_{i}=\delta_{i,2}$), showing decreasing $t_{\text{co}}$ for increasing $w$ within the Mott regime (purple).
@@
The minimal chain to resolve the spin renormalization consists of four quantum dots modeling the central Hubbard system. Thereby, two Hubbard dimers are coupled by the tunable tunneling $w$ (see Fig.7 (a) and (b)). Then, depending on $w$, we probe two regimes: first for $v>w$ (Fig.7 (a)), the system behaves as two uncoupled Hubbard dimers. In the strongly correlated limit its groundstate is then effectively described by two spin-singlets along the outer bonds. Secondly, for $w>v$ (Fig.7 (b)), the situation is quite different, now it forms a spin-singlet at the central bond and between the spins a the edges. As a result, the spin correlation along the probed, lower bond $\langle\overrightarrow{S}_{1}\cdot\overrightarrow{S}_{2}\rangle$ is maximal in the first case ($v>w$) since both the spins at site 1 and site 2 are maximally entangled. Whereas in the second case ($v<w$), $\langle\overrightarrow{S}_{1}\cdot\overrightarrow{S}_{2}\rangle$ is decreased, as both the spins at site 1 and site 2 are essentially uncorrelated. Via Eq.(55), the correlation function significantly alters the hopping in the GFZ Hamiltonian $H^{z}$ at the probed bond and eventually the measured co-tunneling amplitude $t_{\text{co}}$. In Fig.7 (c) we show the expected co-tunneling amplitude as function of the central tunneling $w$ for various correlation strengths $U$. When the system is deep in the Mott regime (large $U$, dark purple), the tunneling amplitude remains nearly constant for $v<w$ but decreases for $w>v$. Remarkably, this effect persists even outside the optimal Mott regime and only vanishes beyond the Mott transition. Where in the weakly interacting regime (green), the trend is slightly inverted: the co-tunneling amplitude increases with larger $w$, ruling out alternative explanations for the tunneling suppression observed in the Mott regime. In Fig.7 (d) we normalized each curve by the initial value at $w=0$, to highlight this effect . For an experimental verification, it is sufficient to observe the decreasing co-tunneling amplitude with increasing $w$, providing a clear qualitative signature of the effect. Precise fine-tuning or knowledge of the tunnelings $v,w$, correlation strength $U$ or the lead-system coupling $C$ is not necessary.

### v2

@@ A.6 Self-consistent selfenergy construction
Assuming a small $\mathcal{S}^{-1}\tilde{G}^{-1}G_{0}^{-1}||<1$ (self-consistency condition), we can expand the right site as a Neumann series and find:
@@
As last step we multiply by $\mathcal{S}^{-1}$ from right and expand both $\mathcal{G}_{\text{H}}$ and $\mathcal{S}^{-1}$, in orders of $\phi$ (up to the highest known order $n$ of $\mathcal{G}_{\text{H}}$ ):
[…]
@@
As an minimal example we determine the selfenergy corresponding to the previously derived GF $G$ in Eq. (34):
@@
a natural decomposition is:
$\tilde{G}^{-1}=-\left[H^{z}-\omega\right]\quad\text{ and }\quad\mathcal{G}_{\text{H}}=-\frac{\phi^{2}}{D^{2}}\mathbb{I}\text{.}$
@@
$\mathcal{S}^{-1}_{0}=0\quad\mathcal{S}^{-1}_{1}=0\quad\mathcal{S}^{-1}_{2}=\frac{1}{D^{2}}\mathbb{I}\text{.}$
@@
$\Sigma$ $\approx\frac{D^{2}}{\phi^{2}}\left[\omega-H^{z}\right]^{-1}\text{,}$ (95)
what is inline with [33].
A.7 Convergence of the Green’s Function Zero Dispersion
Our low-energy GF-approximation is valid in the strongly correlated limit. To probe the convergence of the GFZ-dispersion against our analytical prediction, we computed the exact GFZ s using ED for increasing interaction strengths $U$. In Fig. 5, we present the exact frequency of the GFZ for a fixed momentum $k=\pi$. As can be clearly seen, the exact results converge rapidly to our prediction with increasing correlation strength, as expected. Surprisingly, even at a relatively weak interaction strength of $U=5$, the absolute error for the given parameters is already of order $\sim 10^{-1}$. In the inset, we show the mean distance between the two dispersions and find, on a double-logarithmic scale, the expected $\sim 1/U^{2}$ scaling of the absolute error.
Figure 5: Convergence of the GFZ-dispersion $\varepsilon_{\text{ED}}$ (black) against our analytical prediction in the strongly correlated limit $\varepsilon_{\text{ana}}$(red), exemplarily for $k=\pi$. SSH-Hubbard chain ($N=8$, $w=v=1.0$, PBC, ED). Inset: distance $\Delta_{\varepsilon}=|\varepsilon_{\text{ED}}-\varepsilon_{\text{ana}}|$ and fit (dashed).
Appendix B Perturbation theory for two additional lead quantum dots
@@
$C$ $=\sum_{j,\sigma}l_{j}c^{\dagger}_{j\sigma}c_{L,\sigma}+r_{j}c^{\dagger}_{j\sigma}c_{R,\sigma}+\text{h.c.},$ (97)
@@
$\mathcal{H}^{D}$ $=\{|0,\uparrow\rangle,|0,\downarrow\rangle,|\uparrow,0\rangle,|\downarrow,0\rangle\}$ (100)
@@
The GFZ-Hamiltonian $H^{\text{z}}$ was already derived in Eq. (55), while $H_{0}$ denotes the bare non-interacting Hamiltonian. The result is valid for $\omega\ll U$, where either $\alpha=0$ or $|\alpha|\geq 2$. For $0<|\alpha|<2$ the expansion is close to the singularities in the Hubbard band. Together with Eq. (102) we can further simplify:
@@ Appendix C Charge stability diagram
We compute the stability diagram by determining the global groundstate of the full Hamiltonian $H_{\text{tot}}$ in Eq. (96) for varying weak dot potentials $\mu_{\text{L}},\mu_{\text{R}}$. The groundstate obeys a half-filled Hubbard system for small potentials $\mu_{\text{L}},\mu_{\text{R}}$, while there are either zero $(0,0)$, one $(1,0)$ (dominantly at the left dot), $(0,1)$ (dominantly at the right dot) or two $(1,1)$ particles on the auxiliary dots. We then determine the dot occupations $N_{L}(\mu_{\text{L}},\mu_{\text{R}}),N_{R}(\mu_{\text{L}},\mu_{\text{R}})$. With the onset of the co-tunneling induced coupling between the auxiliary dots, we find that there is a smooth transition between the two localized regimes $(1,0)$ and $(0,1)$. The resulting charge curve $N_{L}(\mu_{\text{L}},\mu_{\text{R}})$ $N_{R}(\mu_{\text{L}},\mu_{\text{R}})$ in this regime can be derived from Eq. (110). Let us rewrite:
@@
Where we used that $a_{s}(\omega)$ and $b_{s}(\omega)$ are linear in the frequency $\omega$. Now for the eigenstates we can neglect the overall shift $C_{0}$ and for a fixed $\bar{\mu}$ we introduce the shifted potential difference $\tilde{\delta}\mu=\delta\mu+\frac{C_{1}\bar{\mu}+C_{3}}{C_{2}}$:
@@
$|\varepsilon_{-}\rangle$ $=\frac{1}{\sqrt{\left(\chi-\sqrt{\chi^{2}+1}\right)^{2}+1}}\begin{pmatrix}\chi-\sqrt{\chi^{2}+1}\\ \frac{t_{\text{co}}(\bar{\mu})}{|t_{\text{co}}(\bar{\mu})|}\end{pmatrix}$ (118)
@@
The occupations on the dots $N_{L}(\bar{\mu},\tilde{\delta}\mu)$$N_{R}(\bar{\mu},\tilde{\delta}\mu)$ are then given by:
@@
The Full-Width-at-Half-Maximum is then given by $\Delta_{FWHM}=2\sqrt{2^{\frac{2}{3}}-1}\frac{|t_{\text{co}}(\bar{\mu})|}{C_{2}}\approx 1.53\frac{|t_{\text{co}}(\bar{\mu})|}{C_{2}}$. Moreover, we can approximate $C_{2}\approx\frac{1}{2}+\mathcal{O}(\frac{\gamma^{2}}{U^{2}})$ and in the case of a strictly local system-dot couplings $l_{k}\propto\delta_{ik},r_{k^{\prime}}\propto\delta_{jk^{\prime}},i\neq j$, the co-tunneling $t_{\text{co}}$ is also independent of $\bar{\mu}$. In total, at zero temperature, we find a universal curve with $\Delta_{FWHM}\approx 3.07|t_{\text{co}}|$.
Appendix D Deviations beyond our perturbative regime for state-of-the-art quantum dot parameters
Figure 6: Co-tunneling $t_{\text{co}}$: numerical results (ED) vs analytical results (Eq. (113)) for the SSH-Hubbard chain ($N=16$, $w=0.8v$, PBC, $l_{i}=\delta_{i,1}$$r_{i}=\delta_{i,2}$). In (a): Co-tunneling $t_{\text{co}}$ as function of $\alpha$ and different relative interaction strengths $\phi$, $D=100$ fixed, numerical results(dashed, green) and analytical results(solid,red). In (b): relative error of $t_{\text{co}}$ determined by Eq. (113) compared to the exact result for various $D$,$U$ and fixed $\gamma=30$. In (c): absolute value of $t_{\text{co}}$, same parameters as in (b). As reference: contour lines of constant $\alpha$, $\phi$ and $t_{\text{co}}$ (grey, green,black).
@@ Appendix D Deviations beyond our perturbative regime for state-of-the-art quantum dot parameters
Both are sufficient to describe the universal behavior, as a global rescaling of the original parameters does not affect the underlying physics. The previously discussed parameter $\phi$ characterizes the Mottness of the central Hubbard system, while $\alpha$ represents the ratio of $\gamma$ to the well-known exchange coupling $J=\frac{D^{2}}{U}$, which is related to the half-bandwidth of the effective low-energy Heisenberg model. In Fig. 6(a), we present the co-tunneling amplitude $t_{\text{co}}$ as function of the coupling ratio $\alpha$ and for various relative interaction strengths $\phi$ in the SSH-Hubbard chain ($w=0.8v$, PBC, $D=(v+w)/2$) with local couplings $l_{i}=\delta_{i,1}$ and $r_{i}=\delta_{i,2}$:
@@
We find that the exact results (dashed, green) show excellent agreement with our analytical approach (solid, red), as given by Eq. (110), already at $\phi\approx 0.2$ and up to relatively large values of $\alpha\approx 2$. As expected, increasing Mottness and weaken the lead-system coupling cause the exact results to converge toward the analytical predictions. More surprising, however, is the large value $\alpha$ with a good agreement, which we interpret as effect of the vanishing first-order contributions of the weak lead-system hopping $C$ in Eq. (98). The drastic change for $\phi>0.2$ signals the Mott transition. From Fig. 6(a), the bare parameters $\gamma,U$ and $D$ can be directly extracted. For example, for $\phi=0.2$ and $\alpha=1.5$ let us fix a half-bandwidth $D=100\mu eV$, then we obtain $U=500\mu eV$, $\gamma=30\mu eV$, and $t_{\text{co}}\approx 1\mu eV$.
@@
Appendix E Renormalization of the GFZ-Hamiltonian by the spin-spin correlation functions
@@
Figure 7: Staggering for OBC inverting at the boundary close to topological transition $v=w$. Here:ED,$U=200$ $N=20$, $\gamma=2$, $w=20$, OBC), with $l_{i}=\delta_{i,1}$$r_{i}=\delta_{i,2}$ (violet) and $l_{i}=\delta_{i,3}$$r_{i}=\delta_{i,4}$ (green).
@@
We comment on the renormalization of the GFZ-Hamiltonian of Eq. (55) by the spin-spin correlation functions. In particular, we focus on the topological features of the underlying SSH model. We consider a finite-size SSH-Hubbard chain, where the non-interacting system shows zero-energy edge modes as a function of the staggered hopping ratio $v/w$. As described by to Eq. (55), the bare hoppings get renormalized significantly by the spin-correlation values. Since the SSH model at large U is effectively a spin chain, its ground state will obey Affleck-Kennedy-Lieb-Tasaki (AKLT) physics [53], with the formation of spin-singlet dimers in correspondence with the stronger bonds. As a consequence, the original imbalance between the non-interacting $v$ and $w$ is further amplified. In Fig.(7) we show the behavior of the renormalized ratio $\frac{\tilde{v}}{\tilde{w}}$ as a function of $v$, for $w=20$ and $U=200$. The grey line shows the behavior of the bare hopping ratio. It is immediate to notice how the renormalized ratio varies much more heavily, passing from strong suppression to strong enhancement. The purple and green curves in the plot refer to different positions along the finite SSH chain, respectively near the boundary and deep in the bulk. Finite-size effect entail a different renormalization in the two cases, and the topological transition, denoted by the red dots, happens at the boundary sites for a lower value of $v/w$ than in the bulk. In the region between the two red dots, the boundary sites are already in the trivial phase, while the bulk is topological. As a consequence, edge state reconstruction ensues [54], with the topological edge zero described in [33] shifting into the bulk.
Appendix F Minimal chain to resolve the renormalization
@@
Figure 8: In (a),(b),(c): small chain of six quantum dots (four as Hubbard system, two as leads). The lead dots are at site 1 and 2 with $l_{i}=\delta_{i,1}$$r_{i}=\delta_{i,2}$. In (a): Two weakly coupled dimers, groundstate approximately described by spin-singlets at bonds 1-2 and 3-4, the probed $t_{\text{co}}$ at bond 1-2 is enhanced by the spin correlation $\langle\overrightarrow{S}_{1}\cdot\overrightarrow{S}_{2}\rangle$. In (b): groundstate resembles spin-singlet at bonds 2-3 and 1-4, spin at site 1 and 2 are roughly uncorrelated and probed $t_{\text{co}}$ at bond 1-2 is suppressed. In (c): large coupling at bond 3-4 restoring initial singlet structure, spins at site 1 and 2 are correlated again and $t_{\text{co}}$ is enhanced. In (d): parameter trace for which we computed the expected $t_{\text{co}}$, passing through all three regimes. In (e) and (f): expected (normalized) co-tunneling $t_{\text{co}}$ as function of $s$ and $U$ (ED, $N=12$, fixed $U=400$ on the lead dots, $v_{1}=100$, $l_{i}=\delta_{i,1}$$r_{i}=\delta_{i,2}$, $\gamma=20$).
@@
The minimal chain to resolve the the renormalization by the spin-spin correlators consists of four quantum dots modeling the central Hubbard system. Thereby, they are coupled pairwise by the tunable tunnelings $w,v_{1},v_{2}$ (see Fig.8 (a), (b), (c)). Then, by varying $w$ and $v_{2}$ ($v_{1}$ is fixed), we probe three regimes: first for $v_{1},v_{2}>w$ (Fig.8 (a)), the system behaves as two uncoupled Hubbard dimers. In the strongly correlated limit its groundstate is then approximately described by two spin-singlets along the outer bonds (1-2, 3-4). Secondly, for $w>v_{1},v_{2}$ (Fig.8 (b)), the situation is quite different, now we find to a good approximation a spin-singlet at the central bond (2-3) and a singlet between the spins at the edges (1-4). In the last case (Fig.8 (b)), $v_{2}>w>v_{1}$, there are again two spin-singlets along the outer bonds (1-2, 3-4) as in (a). As a result, the spin correlation along the probed, lower bond $\langle\overrightarrow{S}_{1}\cdot\overrightarrow{S}_{2}\rangle$ strongly varies for all three regimes. In the first case (a), both the spins at site 1 and site 2 are almost maximally entangled, the probed spin correlation $\langle\overrightarrow{S}_{1}\cdot\overrightarrow{S}_{2}\rangle$ becomes large. Whereas in the second case (b), $\langle\overrightarrow{S}_{1}\cdot\overrightarrow{S}_{2}\rangle$ is decreased, as both the spins at site 1 and site 2 are essentially uncorrelated. Furthermore, deep in the regime (c), the initial singlet pairs are restored and the spin correlation gets large again. As we found in the letter, via Eq. (55) the correlation function significantly alters the hopping in the GFZ Hamiltonian $H^{z}$ at the probed bond and in consequence the measured co-tunneling amplitude $t_{\text{co}}$. Remarkably, in the presented setup Fig.8, it is possible to tune the probed co-tunneling amplitude at bond 1-2 only by varying the couplings $w,v_{2}$ along the other bonds, while $v_{1}$ is fixed. This offers a way to probe the renormalization by the spin-spin correlation functions of the GFZ in an actual experiment by a clear qualitative signature. To illustrate this, we computed the expected co-tunneling amplitude $t_{\text{co}}$ in Fig.8 (e) for all three regimes and different correlation strengths $U$, by probing a trace of various coupling parameters $w,v_{1},v_{2}$ as shown in Fig.8 (d). As predicted by Eq. (55), for the regime (a) ($v_{1},v_{2}>w$) the co-tunneling amplitude $t_{\text{co}}$ is large and we find an almost constant plateau for $s>0$ in Fig.8 (e). By entering the regime (b) ($w>v_{1},v_{2}$) the singlet structure changes, resembling the situation Fig.8 (b). The spins at site 1 and site 2 are less correlated and $t_{\text{co}}$ is suppressed. In Fig.8 (e) for $0.25<s<0.75$ this leads to an drop of $t_{\text{co}}$, which is more pronounced the larger $w$ is compared to $v_{1},v_{2}$. It is now possible to reach again the initial co-tunneling amplitude in the regime (a) by increasing the coupling $v_{2}$ at the bond 3-4. In regime (c) ($v_{2}>w>v_{1}$), the initial singlet pairings are approximately restored and the co-tunneling $t_{\text{co}}$ is again enhanced as we see in Fig.8 (e) for $s>0.75$. Surprisingly, the described effect at bond 1-2 is reached by tuning only the couplings at the 2-3 and 3-4 bonds while $v_{1}$ at 1-2 is always constant (see Fig.8 (d)). However, the observation becomes clear by understanding how the spins are roughly paired in the groundstate and how this affects the probed GFZ. In Fig.8 (f), we normalized each curve by the initial value at $s=0$, to highlight this effect. Our presented setup offers a qualitative signature, precise fine-tuning or knowledge of the tunnelings $v_{1},v_{2},w$, correlation strength $U$ or the lead-system coupling $C$ is not necessary as long as one probes roughly in our perturbative regime $\gamma\ll w,v_{1},v_{2}$, $w,v_{1},v_{2}<U$.